Matthew K. Gray

Work Address MIT Room E15-468a 20 Ames Street Cambridge, MA 02139

(617) 253–1750 Home Address 64 Webster Ave Cambridge, MA 02141 (617)-497–3934 mkgray@mit.edu

Experience  MIT Media Lab, Personal Information Architecture  Cambridge, MA

May, 1997 – present Researched networking infrastructure for, and applications of statistical inference systems to “Things That Think”, a large scale project on ubiquitous computing in everyday objects. Responsibilities include those described below, as well as supervising undergraduates, team management, and high level project design.

MIT Media Lab, Physics and Media Cambridge, MA April, 1996 – May, 1997 Research into electrostatic intrabody signaling. Responsibilities include hardware and software design, implementation, testing, presentation, and handling sponsor relationships.

net.Genesis Corporation Cambridge, MA April, 1994 – January, 1996 Founder and Chief Technologist for a company offering server side software tools for web developers. Developed initial versions of products, did product design and strategic planning. Other responsibilities included coding, sales support, marketing support, customer service, and management of engineering teams.

O’Reilly and Associates Cambridge, MA May, 1994 – August, 1994 Consulting work focusing on advertiser reports and web server enhancements.

Extensive prior work in undergraduate physics research at MIT and Brookhaven National Labs. Work included data collection, analysis and presentation.

Education  Massachusetts Institute of Technology  Cambridge, MA

M.S. in Media Technology, May 1999

Massachusetts Institute of Technology Cambridge, MA B.S. in Physics, May 1997

Publications, Organizations, Proficiencies I wrote a book Build a Web Site, Prima Publishing, June, 1995, covering technical and design aspects of managing web sites. I have also written extensively on the growth of the web, based on research started in 1993. This research has been cited in Newsweek, Scientific American, PC Week, PC World and many others.

I am a member and former chairman of the Student Information Processing Board, a volunteer organization providing computing services to the MIT community. I set up one of world’s first 100 web servers in May 1993 as part of this group.

I am a member of the Apache Group, a volunteer development group that developed the Apache Web Server, the most popular web server in the world.

I am fluent in a wide range of computer languages, C, perl, Java, and many others.